OpenStudy (anonymous):

Write an equation for a linear function f that has the given value: f(-2)=-2, f(4)=-8

OpenStudy (anonymous):

y=-x-4. shall i simplify??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(x)=mx+n f(-2)=-2m+n=-2 f(4)=4m+n=-8 (4m+n)-(-2m+n)=-8-(-2) 4m+n+2m-n=-6 6m=-6 m=-1 4m+n=-8 4*(-1)+n=-8 -4+n=-8 n=-4 f(x)=-x-4 did you get it??

OpenStudy (anonymous):

oh its simple lol. Thank you.

OpenStudy (anonymous):

you're welcome :)